A swimming pool is 50metres long and 15metres wide.Its shallow and deep ends are 3/2 metres and9/2 metres deep respectively.If the bottom of the pool slopes uniformly, find the amount of water required to fill the pool.
Answer by rothauserc(4718)

1) find the average depth of the pool by dividing the sum of the depths of the shallow end and deep end by 2
3/2 + 9/2 = 12/2 = 6
average depth = 6 / 2 = 3
2) volume of pool is length * width * average depth
50 * 15 * 3 = 2250 cubic metres of water are necessary to fill the pool